OpenStudy (anonymous):

Given that f(x)=4/x-3,x not equal to 3 find the value of x if 3f(x)=2/x , x not equal to 0

OpenStudy (anonymous):

3f(x)=2/x 3(4/x-3)=2/x 1/x(12-2)=9 x=10/9